sensitive as 0.55

specificity as 0.45

Questions ...

1-what is the positive predictive value of the digit rectal exam for prostate cancer in men if the prevalence is 21 per 1,000 men .

2- what is the probability of a false negative test

3- what is the probability of a false positive test
Answer #1

1)

P(tested positive) =P(have prostate cancer)*P(tested positive|have prostate cancer)+P(not have prostate cancer)*P(tested positive|not have prostate cancer)

=(21/1000)*0.55+(1-21/1000)*(1-0.45)=0.55

positive predictive value

=P(have prostate cancer)*P(tested positive|have prostate cancer)/P(tested positive)

=(21/1000)*0.55/0.55

=0.021

2)

P(false negative) =1- sensitivity =1-0.55 =0.45

3)

probability of a false positive test =1- specificity =1-0.45 =0.55